A stone is projected at anangle of 60 deegree from the ground level strikes at an angle 30 deegree on the roof of a building of height ‘h’.then the speed of projection of the stone is:
(A)2gh0.5  (B)6gh0.5   (C)3gh0.5 (D)gh0.5
 pls give ans with solution....

Grade:11

1 Answers

Vikas TU
7 years ago
Components in x and y direction of initial velocity with 60 degree are:
ucos60 and usin60 respectively.
At h height distance,
tan30 = Vy/ucos60
Vy = u/2root3
Further Distance in y – direction :
Vy^2 – (usin60)^2 = 2as
(u/2root3)^2 – (uroot3/2)^2  = -2gs
On solving we get,
u = root(3gh)
that is (C) option.